Use the 5 steps for problem solving to solve this problem..
Tawny found $1.95 in quarters and dimes in the couch. There were 4 times as many dimes as there were quarters. How many dimes and quarters did she find?
This sounds like a joke instead of a math problem...but its not funny for me Im lost can someone please help me????Thank you very much...

Answer by htmentor(1343)

Tawny found $1.95 in quarters and dimes in the couch. There were 4 times as many dimes as there were quarters. How many dimes and quarters did she find?
===================
Let q = the number of quarters
Then the number of dimes = 4q [4 times as many dimes as quarters]
The total amount can be written 25q + 10(4q) = 195 [in cents]
Solve for q:
25q + 40q = 195
65q = 195
q = 3
Therefore, the number of dimes = 4*3 = 12
Check:
25*3 + 12*10 = 75 + 120 = 195